OpenStudy (anonymous):

Using Substitution Rule, find: 1) ∫(sec^2(1/x))/x^2 2) ∫(sec^2)2Θ

OpenStudy (lgbasallote):

\[\huge \int \frac{\sec^2 \left(\frac 1x \right)}{x^2} dx\] let \[u = \frac 1x\] \[du = -\frac 1{x^2}\] so your integral becomes \[\huge \implies -\int \sec^2 u du\] does that help?

OpenStudy (anonymous):

Yes, it does. Thank you! :)

OpenStudy (lgbasallote):

welcome

OpenStudy (lgbasallote):

for the second part \[\huge ]int \sec^2 (2\theta) d\theta\] let \[u = 2\theta\] \[du = 2d \theta\] so your integral becomes \[\huge \implies \frac 12 \int \sec^2 u du\]

OpenStudy (anonymous):

Thank you! x 10000000 :)

OpenStudy (lgbasallote):

welcome